A particle of mass m=5 is moving with uniform speed v=3โˆš2 in the XY plane along the line y=x+4. The magnitude of angular momentum about origin is:


ans: 60


ย 


first of all, how can a body undergoing linear motion have angular momentum.


ย 


secondly, they have used L=mvr and taken r as 4/โˆš2. How?

Grade:11

3 Answers

Fawz Naim
12 years ago

the body is undergoing rotational motion about the origin i.e. it is rotating in a circle having its centre as the origin and the radius is equal to the distance of that point from the origin. the distance of the given line from origin is 4/square root2

now the angular momentum can be calculated easily by putting the values

snehashish bal
12 years ago

see one can calculate da angular momentum undergoing linear motion from da origin......nw u draw da coordinate axes and draw da line u ll find dat da prependicular distance of da origin to the line is 4/V2......

vikas askiitian expert
12 years ago

 y = x+4               (given)

 x-y+4 = 0

perpendicular distance from origen = 4/sqrt2=2sqrt2 m

now angular momentam = mvr                           (r is the perpendicular distance of line of motion of particle                                                                      & point about which we have  to calculate angular momentam)

                                 =5*(3sqrt2)2sqrt2

                                 =5*3*4

                                 =60